f(x)=2x^n+a, if f(2)=26 and f(4)=138, then f(3)is equal to

Grade:12

1 Answers

varun jain
12 years ago

Easy man..64…look

2.(2)^n+a=26

And 2(4)^n+a=138…solve to get a=10 and n=3….hencef(3)=2(3)^3+10=64